OpenStudy (anonymous):

if a ratio is 1:3200 what would x:6,775,235,700 be

OpenStudy (anonymous):

\[\left( 1\over3200 \right)=\left( x \over6 \right)\]

OpenStudy (anonymous):

\[x = \left( 6\over3200 \right) = \left( 1\over1200 \right)\]

OpenStudy (anonymous):

oops. sorry

OpenStudy (anonymous):

i simplified it wrongly. just stick with 6/3200

OpenStudy (anonymous):

the other cases are similar.

OpenStudy (anonymous):

91 rt you see a satillite is going to hit the earth and the estimated human casulties according to nasa is 1 in 3200 my question is when you apply that to the world's population of 6,775,235,700 you mean it would be 6 in 3200?